# GuideGen: A Text-Guided Framework for Full-torso Anatomy and CT Volume Generation

### v2

@@ 6.2.3 Utilizing GuideGen for Dataset Synthesis or Augmentation
GuideGen 0.025 0.041 0.017 0.132 0.147 0.152 0.141 0.084 0.088 0.271 0.152 0.240 - 0.112
Table 11: Segmentation variance on multi-organ segmentation tasks using real or synthetic cases from each method. The significance of our improvement over the second-best framework is tested using a two-sided $t$-test, with a $p$-value of 0.016 and 0.001 respectively for two datasets. Abbreviations: Spl.: Spleen, Kid.: Kidneys, Sto.: Stomach, Pan.: Pancreas, Adr.: Adrenal Glands, Eso.: Esophagus, IVC: Inferior Vena Cava, Gall.: Gallbladder, Duo.: Duodenum, Blad.: Urinary Bladder, PV&SV: Portal Vein and Splenic Vein. Some of the semantic classes reported are not consistent with the original dataset since they are merged as described in 6.1.2.
@@
Figure 5: Samples generated by GuideGen for qualitative assessment. Five slices are shown in each case along with the multi-organ anatomies and tumor semantics (if any). We mark the bounding box of tumor labels in the CT slices shown using red outlines. Note that the first four cases accommodate a tumor site for the clinical description in the medical prompt while the last two has no tumor indication from the input prompt and generate a healthy anatomy and CT instead.
@@
Figure 7: Qualitative results for tumor segmentations on MSD-Lung, MSD-Colon and KiTS datasets. One slice is chosen for each case, which is centered at the largest transversal section of the tumor.
@@
Dataset No. Real cases No. Synthetic cases for training / Avg. DSC
for training 0 100 200 500 1000
BTCV 0 0.00 0.69 0.79 0.79 0.81
24 0.74 0.81 0.82 0.83 0.83
AMOS 0 0.00 0.64 0.79 0.81 0.81
240 0.84 0.82 0.85 0.87 0.87
MSD-Lung 0 0.00 0.64 0.69 0.71 0.73
40 0.69 0.72 0.75 0.77 0.78
MSD-Colon 0 0.00 0.07 0.24 0.27 0.33
80 0.47 0.41 0.49 0.50 0.48
KiTS 0 0.00 0.46 0.51 0.59 0.63
210 0.72 0.72 0.72 0.74 0.77
Table 14: Performance gain by using different numbers of GuideGen’s generated samples for segmentation dataset synthesis or augmentation. We report the average DSCs for segmentation performance on each dataset. For reference, the main article reports the performance of segmentation models trained using all real training cases as opposed to those trained using 200 synthetic cases for each dataset.

### v3

@@ Appendix C Limitation and Future Directions
GuideGen 0.021 0.056 0.021 0.099 0.195 0.130 0.277 0.060 0.108 0.199 0.213 0.297 - 0.140
Table 10: Segmentation variance on multi-organ segmentation tasks using real or synthetic cases from each method. The significance of our improvement over the second-best framework is tested using a two-sided $t$-test, with a $p$-value of 0.016 and 0.001 respectively for two datasets. Abbreviations: Spl.: Spleen, Kid.: Kidneys, Sto.: Stomach, Pan.: Pancreas, Adr.: Adrenal Glands, Eso.: Esophagus, IVC: Inferior Vena Cava, Gall.: Gallbladder, Duo.: Duodenum, Blad.: Urinary Bladder, PV&SV: Portal Vein and Splenic Vein. Some of the semantic classes reported are not consistent with the original dataset since they are merged as described in A.2.
@@
Figure 6: Masks and Images generated by GuideGen for qualitative assessment. 6 slices are shown in each case along with the multi-organ anatomies and tumor semantics (if any). We show the input medical prompts on the left and mark the tumor labels in the CT slices shown in red.
@@
Figure 8: Qualitative results for tumor segmentations on LU, CO and KI datasets. One slice is chosen for each case, which is centered at the largest transversal section of the tumor.
@@
[…]
LU CO KI Avg. LU CO KI Avg.
GuideGen 0.71 0.21 0.64 0.52 8.4 227.0 84.5 106.6
DiffTumor (chen2024towards) 0.55 0.12 0.49 0.39 46.4 291.5 92.2 143.4
LesionDiffusion (tian2025lesiondiffusion) 0.69 0.36 0.63 0.56 11.0 128.3 79.1 72.8
Table 13: Performance of tumor segmentation model pretraining compared to tumor inpainting methods. We report the DSC and HD95 metrics of three tumor segmentation datasets as in the main paper.